1) The one-sample t-statistic for testing H0: μ = 0 Ha: μ > 0 from a sample of n=20 is t=1.84

a. What are the degrees of freedom for this statistic?

b. What are the two critical values of t* that bracket t= 1.84 from the t-table?

c. Is the value t=1.84 significant at both the 5% and 1% level?
Answer #1

a)

Degree of freedom: df=n-1=20-1=19

b)

The two critical values for df=19 bracket t=1.84 are 1.729 and 2.093.

c)

The p-value is between 0.025 and 0.05. That is

0.025 < p-value < 0.05

Since p-value is less than 0.05 so it is significant at 5% level of significance.

Since p-value is not less than 0.01 so it is not significant at 1% level of significance.